OpenStudy (anonymous):

X^2 - 7x - 18 = ? Medal will be given, help?

OpenStudy (anonymous):

\[x^{2}-7x-18=0\]\[(x-9)(x+2)=0\] Case 1: \[x-9=0\]\[x=9\] Case 2:\[x+2=0\]\[x=-2\]

OpenStudy (anonymous):

You can check with a plot.

OpenStudy (anonymous):

Thank you so much! :)